Find all solutions to the equation x+sqrt(x-2)=4
Answer by lwsshak3(11628)

Find all solutions to the equation
x+sqrt(x-2)=4
√(x-2)=4-x
square both sides
x-2=16-8x+x^2
x^2-9x+18=0
(x-3)(x-6)=0
x=6(extraneous root)
or
x=3